Suppose two buckets contain sunflower seeds at a ratio of 7:3. If the bucket with fewer seeds contains 168 seeds, how many seeds are in the other bucket?
Answer by Theo(13342)

7/3 = x/168
cross multiply to get:
7*168 = 3*x
divide both sides of the equation by 3 to get:
7*168/3 = x
solve for x to get:
x = 392
7/3 = 392/168
cross multiply to get:
7 * 168 = 3 * 392
simplify to get:
1176 = 1176
this proves the ratios are equal.
your solution is that the bucket with more seed contains 392.